Instructions: Perform indicated operations.
What is the solution for (x-4(x+7)

Answer by 303795(562)

x - 4(x + 7) Multiply the -4 by each of the items inside the bracket( -4 *x = -4x and -4 * 7 = -28)
= x - 4x - 28 Add like terms (x - 4x = -3x)
=-3x - 28